# Heat dissipation in marginally stable linear time-delayed Langevin systems

### v3

@@ Abstract
The thermodynamic properties of time-delayed dynamics remain largely unexplored, especially for systems that exhibit asymptotically non-stationary behavior. Here, we investigate heat dissipation in two classes of marginally stable linear time-delayed Langevin dynamics: (i) diffusive criticality, which asymptotically manifests as scaled Brownian diffusion, and (ii) oscillatory criticality, which shows oscillation with diffusive amplitude. By analytical derivations, we find fundamentally different thermodynamic signatures: the average heat dissipation rate asymptotically approaches a constant for diffusive criticality but diverges linearly with oscillations for oscillatory criticality, despite both showing linearly growing variance over time. We further numerically study the probability distributions of heat dissipation rates for both types of critical dynamics. Our results demonstrate that non-stationary time-delayed dynamics with similar scaling can yield qualitatively distinct heat dissipation behaviors, depending on the underlying dynamical details. This work provides a concrete foundation for future investigations into thermodynamic properties of general nonlinear time-delayed systems.
@@ Introduction
Time-delayed systems are ubiquitous in natural and engineered systems [1, 2, 3, 4, 5, 6, 7], where finite signal propagation or processing time introduces nontrivial memory effects. Time-delayed effects may induce complex phenomena, including spontaneous oscillations [8, 7, 9], bifurcations [10, 11], resonance [12, 13, 14], chaos [15], synchronization [16, 17, 18], and pattern formation [19, 20]. Investigating these delay-induced effects is crucial for the accurate analysis and effective design of real-world systems.
@@
Time-delayed systems can also manifest nontrivial thermodynamic properties due to their inherent non-Markovian nature. Several studies reveal non-zero (sometimes even negative) average heat dissipation in time-delayed Langevin systems in the steady state, highlighting their non-equilibrium characters [21, 22, 23, 24, 25, 26]. Thermodynamic quantities are also discussed using generalized fluctuation theorems and thermodynamic uncertainty relations adapted for time-delayed dynamics [27, 28, 29, 30, 31, 32, 33, 34]. However, current research on the thermodynamic properties of time-delayed systems has primarily focused on stable dynamics (i.e., dynamics that permits a steady state). For systems in non-stationary states, including those exhibiting diffusive or oscillatory behavior under Pyragas time-delayed feedback control [35, 36, 37], the thermodynamic properties have not been comprehensively explored. Even for linear time-delayed Langevin dynamics, a systematic study is lacking on the heat dissipation at marginal stability (when the parameters of the system lie on the boundary of the stability conditions [38]).
@@
To address this gap, we study the heat dissipation in two classes of marginally stable linear time-delayed Langevin dynamics: (i) diffusive criticality and (ii) oscillatory criticality. We evaluate the average heat dissipation rate through the variance of the system’s position, which is asymptotically determined by the roots of the linear stability equation with zero real parts. Surprisingly, we find that although both types of criticality exhibit linearly divergent variance in time, they show fundamentally different thermodynamic behaviors: the average heat dissipation rate approaches a constant in the long-time limit for diffusive criticality but diverges linearly in time for oscillatory criticality. Moreover, we also examine the distribution of the heat dissipation via numerical simulations, which further reveal distinct statistical properties for each criticality class. Our results for these linear systems demonstrate that non-stationary time-delayed dynamics with similar scaling can produce qualitatively different heat dissipation, depending on the details of the dynamics. This study provides a concrete example describing the thermodynamic properties in time-delayed Langevin systems beyond the stable regime, laying a foundation for broader investigations into nonlinear systems.
@@
This work is structured as follows: Sec. II introduces the linear time-delayed Langevin dynamics and specifies the two classes of marginally stable dynamics. Sec. III presents the core of our work: the analytical derivation of the asymptotic average heat dissipation rates for both critical dynamics, alongside numerical results that validate our derivation and explore the full distribution of dissipation. Finally, Sec. IV summarizes this work and discusses the future directions.
@@
where $\tau>0$ is a fixed delay time, $a$ denotes the strength of a linear non-delayed force, and $b$ represents the strength of the linear time-delayed force. The friction coefficient is given by $\gamma$, and $\xi(t)$ is white Gaussian noise satisfying $\left\langle\xi(t)\right\rangle=0$ and $\left\langle\xi(t)\xi(t^{\prime})\right\rangle=2\gamma T\delta(t-t^{\prime})$, where $T$ is the temperature of the thermal environment. The dynamics is initialized at $t=t_{\text{init}}$ subjected to a predefined historical function $\phi(t)$ defined for $t\in[t_{\text{init}}-\tau,t_{\text{init}}]$. Throughout this work, we set $t_{\text{init}}=0$ and $\gamma=1$ without loss of generality.
The solution to the dynamics described by Eq. (1) is as follows [38]:
$x(t)$ $=x_{0}(t)\phi(0)+b\int_{-\tau}^{0}x_{0}(t-s-\tau)\phi(s)ds$
$\quad+\int_{0}^{t}x_{0}(t-s)\xi(s)ds,$ (2)
@@
In Appendix B, we show that
$x_{0}(t)=\sum_{k=-\infty}^{\infty}\dfrac{1}{1+b\tau e^{-S_{k}\tau}}e^{S_{k}t}\ \ (t>0),$ (4)
@@
$\langle x(t)\rangle$ $=x_{0}(t)\phi(0)+b\int_{-\tau}^{0}x_{0}(t-s-\tau)\phi(s)ds,$ (6)
$\langle x^{2}(t)\rangle$ $-\langle x(t)\rangle^{2}=2T\int_{0}^{t}x_{0}^{2}(s)ds.$ (7)
Parameters $a,b,\tau$ in Eq. (5) determine the distribution of characteristic roots $\{S_{k}\}$, which further control the asymptotic behavior of the solution $x(t)$ . If $\max\{\text{Re}[S_{k}]\}<0$, the distribution of $x(t)$ becomes stationary as $t\rightarrow\infty$ and the dynamics is stable; if $\max\{\text{Re}[S_{k}]\}=0$, the variance of $x(t)$ diverges linearly in time and the dynamics is marginally stable; if $\max\{\text{Re}[S_{k}]\}>0$, the dynamics exponentially diverges in time.
@@
Figure 1: (a) A single trajectory simulation of the dynamics with diffusive criticality. Here a=5a=-5, $b=5$, $\tau=1$, $T=1$, and $\phi(t)=0\ \ (t\leq 0)$. (b) 100 trajectories of the dynamics with diffusive criticality for each set of parameters. Blue: $a=0.5$ and b=0.5b=-0.5; red: $a=b=0$ (Brownian motion) ; green: a=1a=-1 and $b=1$. Here $T=1$, $\tau=1$, and $\phi(t)=0\ \ (t\leq 0)$.
@@
$a+b=0,\ a\tau<1,$ (8)
there is exactly one zero root in $\{S_{k}\}$ and all the others have negative real parts (we exclude $a\tau=1$ as the coefficients in Eq. (4) diverge). Such critical dynamics corresponds to scenarios of systems under Pyragas control [43, 35, 36, 37]. Under this condition (Eq. (8)), the fundamental solution satisfies
$\lim_{t\rightarrow{\infty}}x_{0}(t)=\dfrac{1}{1-a\tau}.$ (9)
Combining Eq. (9) with Eqs. (6) and (7), we obtain the dynamics in the long-time limit ($t\rightarrow\infty$)
$\langle x(t)\rangle=\frac{1}{1-a\tau}\phi(0)-\frac{a}{1-a\tau}\int_{-\tau}^{0}\phi(s)ds,\ \ (t\rightarrow\infty)$ (10)
$\langle x^{2}(t)\rangle-\langle x(t)\rangle^{2}=\frac{2T}{(1-a\tau)^{2}}\cdot t+C_{0},\ \ (t\rightarrow\infty)$ (11)
@@
$C_{0}=2T\int_{0}^{\infty}\left[x_{0}^{2}(s)-\left(\dfrac{1}{1-a\tau}\right)^{2}\right]ds,$ (12)
is a bounded constant  [^1] The convergence can be verified by noticing $x_{0}(0)=\sum_{k=-\infty}^{\infty}\dfrac{1}{1+b\tau e^{-S_{k}\tau}}=1$ and accordingly $x_{0}(t)=\frac{1}{1-a\tau}+O(e^{\gamma t})$, where $\gamma$ is the largest nonzero real part of the characteristic roots. Similar discussion applies for Eq. (18). See also [53].. Note that Eq. (11) has been analytically derived in Ref. [45]. The mean displacement $\langle x(t)\rangle$ asymptotically approaches a constant that depends on the historical trajectory $\phi(t)$. The diffusion of the critical dynamics resembles free Brownian motion with a scaling factor $1/(1-a\tau)^{2}$ (Fig. 1(a)). A positive time-delayed force (b=a>0b=-a>0) suppresses the diffusion and a negative one ($b=-a<0,\ a\tau<1$) enlarges the diffusion (Fig. 1(b)).
@@
there are exactly two conjugate imaginary roots $\pm i\sqrt{b^{2}-a^{2}}$ and all the others have negative real parts. Such oscillatory critical dynamics shows oscillation with a fluctuating amplitude (Fig. 2(a)). The fundamental solution in the long-time limit is
$\lim_{t\rightarrow\infty}x_{0}(t)$ $=\frac{1}{1+b\tau e^{-i\omega_{c}\tau}}e^{i\omega_{c}t}+\frac{1}{1+b\tau e^{i\omega_{c}\tau}}e^{-i\omega_{c}t}$
$=\dfrac{2\cos(\omega_{c}t+\theta)}{\sqrt{1+b^{2}\tau^{2}-2a\tau}},$ (14)
@@
$\omega_{c}=\sqrt{b^{2}-a^{2}},\ \ \tau=\dfrac{\arccos(-\frac{a}{b})}{\sqrt{b^{2}-a^{2}}},$ (15)
$\theta=\arctan\left(\frac{\sqrt{b^{2}-a^{2}}\cdot\tau}{a\tau-1}\right).$
The mean displacement in the long-time limit ($t\rightarrow\infty$)
$\langle x(t)\rangle=\dfrac{2}{\sqrt{1+b^{2}\tau^{2}-2a\tau}}$
$\cdot\left(\phi(0)\cos(\omega_{c}t+\theta)+b\int_{-\tau}^{0}\cos[\omega_{c}(t-s-\tau)+\theta]\phi(s)ds\right)$ (16)
and the variance
@@
$\dfrac{4T\bigg{[}t+\frac{1}{2\omega_{c}}\bigg{(}\sin[2(\omega_{c}t+\theta)]-\sin(2\theta)\bigg{)}\bigg{]}}{1+b^{2}\tau^{2}-2a\tau}+C_{1},$ (17)
@@
$C_{1}=2T\int_{0}^{\infty}\bigg{(}x_{0}^{2}(s)-\dfrac{4\cos^{2}(\omega_{c}s+\theta)}{1+b^{2}\tau^{2}-2a\tau}\bigg{)}ds,$ (18)
@@
Figure 2: (a) A single trajectory simulation of the dynamics with oscillatory criticality. Here a=2a=-2, b=3b=-3, $\tau\approx 1.029$, $T=1$, and $\phi(t)=0\ \ (t\leq 0)$. (b) 100 trajectories of the dynamics with oscillatory criticality for each set of parameters. Blue: a=2a=-2, b=10b=-10, and $\tau\approx 0.181$; red: a=2a=-2, b=5b=-5, and $\tau\approx 0.433$; green: a=2a=-2, b=2.5b=-2.5, and $\tau\approx 1.665$. Here $T=1$ and $\phi(t)=0\ \ (t\leq 0)$.
@@
In the following, we analytically derive the asymptotic expressions for the average heat dissipation rate of two types of marginally stable dynamics, along with numerical verification. This constitutes the main result of our work.
@@
where $\circ$ denotes the Stratonovich product [46]. It represents the heat dissipated from the system into the surrounding thermal medium per unit time. The heat dissipation rate of steady state for stable solutions has been derived in [21, 22, 26]. However, those methods are not applicable when calculating the heat dissipation of marginally stable dynamics. In this section, we explicitly calculate the heat dissipation for the two classes of marginally stable dynamics.
From the definition in Eq. (19) and the dynamics described by Eq. (1), the heat dissipation of the system within an infinitesimal time interval $dt$ is given by ($\gamma\equiv 1$):
$\langle dq(t)\rangle=\left\langle\bigg{[}\frac{dx(t)}{dt}-\xi(t)\bigg{]}\circ dx(t)\right\rangle$
$=\left\langle\big{[}ax(t)+bx(t-\tau)\big{]}\circ\big{[}ax(t)+bx(t-\tau)+\xi(t)\big{]}dt\right\rangle$
$=a^{2}\langle x^{2}(t)\rangle dt+b^{2}\langle x^{2}(t-\tau)\rangle dt$
$\quad\quad+2ab\langle x(t)x(t-\tau)\rangle dt+aTdt,$ (20)
@@
$\langle x(t)\circ\xi(t)dt\rangle=\left\langle\frac{x(t+dt)+x(t-dt)}{2}\cdot\xi(t)dt\right\rangle$
$=\frac{1}{2}\int_{0}^{t+dt}x_{0}(t-s)\langle\xi(s)ds\cdot\xi(t)dt\rangle$
$=\frac{1}{2}\left(\int_{0}^{t+dt}x_{0}(t-s)\cdot 2T\delta(t-s)ds\right)\cdot dt$
$=\frac{1}{2}\cdot x_{0}(0)\cdot 2Tdt=Tdt\quad(\text{since}\ x_{0}(0)=1)$ (21)
and
$\langle x(t-\tau)\circ\xi(t))=0$ (22)
because the system’s state at past is independent of the white noise at present. The correlation $\langle x(t)x(t-\tau)\rangle$ can be calculated through
@@
$=a\langle x^{2}(t)\rangle dt+b\langle x(t)x(t-\tau)\rangle dt+\langle x(t)\circ\xi(t)dt\rangle,$ (23)
@@
$b\langle x(t)x(t-\tau)\rangle=\frac{1}{2}\frac{d\langle x^{2}(t)\rangle}{dt}-a\langle x^{2}(t)\rangle-T.$ (24)
@@
$\left\langle\frac{dq}{dt}\right\rangle=b^{2}\langle x^{2}(t-\tau)\rangle-a^{2}\langle x^{2}(t)\rangle+a\frac{d\langle x^{2}(t)\rangle}{dt}-aT.$ (25)
Below, we discuss the heat dissipation rate of two types of marginally stable dynamics (we set $\phi(t)=0\ \ (t<0)$ for simplicity).

### v4

@@ Abstract
The thermodynamic properties of time-delayed dynamics remain largely unexplored, especially for systems that exhibit asymptotically non-stationary behavior. Here, we investigate heat dissipation in two classes of marginally stable linear time-delayed Langevin dynamics: (i) diffusive criticality, which asymptotically manifests as scaled Brownian diffusion, and (ii) oscillatory criticality, which shows oscillation with diffusive amplitude. By analytical derivations, we find fundamentally different thermodynamic signatures: the average heat dissipation rate asymptotically approaches a constant for diffusive criticality but diverges linearly with oscillations for oscillatory criticality, despite both showing linearly growing variance over time. We discuss in detail how the heat dissipation rate behaves differently as the dynamics asymptotically approaches these two criticality classes from the stable regime. We also numerically study the probability distributions of heat dissipation rates for both types of critical dynamics. Our results demonstrate that non-stationary time-delayed dynamics with similar scaling of variance can yield qualitatively distinct heat dissipation behaviors, depending on the underlying dynamical details. This work provides a concrete foundation for future investigations into thermodynamic properties of general nonlinear time-delayed systems.
@@ Introduction
Time-delayed systems are ubiquitous in natural and engineered systems [1, 2, 3, 4, 5, 6, 7], where finite signal propagation or processing time introduces nontrivial memory effects. Time-delayed effects may induce complex phenomena, including spontaneous oscillations [8, 7, 9], bifurcations [10, 11], resonance [12, 13, 14], chaos [15], synchronization [16, 17, 18], and pattern formation [19, 20]. Investigating these delay-induced effects is crucial for the accurate analysis and effective design of real-world systems. In the presence of environmental noise, the dynamics of time-delayed systems requires a probabilistic description, such as time-delayed Langevin equations and time-delayed Fokker-Planck equations [21, 22, 23, 24, 25]. For noisy systems, time-delayed effects may induce nontrivial thermodynamic behaviors due to its inherent non-Markovian nature. Several studies have revealed non-zero and even negative average heat dissipation in time-delayed Langevin systems in the steady state, highlighting their non-equilibrium character [26, 27, 28, 29, 30, 31]. Thermodynamic quantities are also discussed using generalized fluctuation theorems and thermodynamic uncertainty relations adapted for time-delayed dynamics [32, 33, 34, 35, 36, 37, 38, 39, 40]. However, current research on the thermodynamic properties of time-delayed systems has primarily focused on stable dynamics (i.e., dynamics that permits a steady state). For systems in non-stationary states, including those exhibiting diffusive or oscillatory behavior under Pyragas time-delayed feedback control [41, 42, 43], the thermodynamic properties have not been comprehensively explored. Even for linear time-delayed Langevin dynamics, a systematic study is lacking on the heat dissipation at marginal stability when the parameters of the system lie on the boundary of the stability conditions [23].
@@
To address this gap, we study the heat dissipation in two classes of marginally stable linear time-delayed Langevin dynamics: (i) diffusive criticality and (ii) oscillatory criticality. We evaluate the average heat dissipation rate through the variance of the system’s position, which is asymptotically determined by the roots of the linear stability equation with zero real parts. Surprisingly, we find that although both types of criticality exhibit linearly divergent variance in time, they show fundamentally different thermodynamic behaviors: the average heat dissipation rate approaches a constant in the long-time limit for diffusive criticality but diverges linearly in time for oscillatory criticality. Asymptotic analysis also reveals distinct behaviors of the heat dissipation rate as the system parameters approach the criticality boundaries from the stable regime. For systems near diffusive criticality, the steady-state heat dissipation rate asymptotically approaches a constant limiting value consistent with the result for diffusive criticality within a finite timescale, despite the linearly divergent variance of the system’s position. For systems near oscillatory criticality, however, the steady-state heat dissipation rate itself diverges asymptotically, requiring an infinitely long time to be attained. Furthermore, the spectral decomposition and distribution of heat dissipation approach a well-defined limiting form near diffusive criticality, but fail to do so near oscillatory criticality. Our results for these linear systems demonstrate that non-stationary time-delayed dynamics with similar scaling of variance can produce qualitatively different heat dissipation, depending on the details of the dynamics. This study provides a concrete example describing the thermodynamic properties in time-delayed Langevin systems beyond the stable regime, laying a foundation for broader investigations into nonlinear systems.
@@
This work is structured as follows: Sec. II introduces the linear time-delayed Langevin dynamics and specifies the two classes of marginally stable dynamics. Sec. III and Sec. IV present the core of this work: Sec. III discusses the analytical derivation of the asymptotic average heat dissipation rates for both critical dynamics, alongside numerical results that validate our derivation and explore the full distribution of dissipation. Sec. IV compares the asymptotic behaviors of the average heat dissipation rate for stable dynamics near the stability boundaries of the two distinct criticality classes. Finally, Sec. V summarizes this work and discusses the future directions.
@@
where $\tau>0$ is a fixed delay time, $a$ denotes the strength of a linear non-delayed force, and $b$ represents the strength of the linear time-delayed force. The friction coefficient is given by $\gamma$, and $\xi(t)$ is thermal noise satisfying $\left\langle\xi(t)\right\rangle=0$ and $\left\langle\xi(t)\xi(t^{\prime})\right\rangle=2\gamma T\delta(t-t^{\prime})$, where $T$ is the temperature of the thermal environment. The dynamics is initialized at $t=t_{\text{init}}$ and the past trajectory in $t_{\text{init}}-\tau,t_{\text{init}}]$ is described by a predefined function $\phi(t)$. Throughout this work, we set $t_{\text{init}}=0$ without loss of generality, and we assume that $\phi(t)$ is an absolutely continuous deterministic function and its derivative is integrable over the interval $\tau,0]$. Physical examples of dynamics Eq. (1) include nanomechanical resonators [33, 44] and colloidal particles [45, 46] under time-delayed feedback control.
According to Ref. [23], the solution to the dynamics in Eq. (1) is as follows:
$x(t)$ $=x_{0}(t)\phi(0)+{\color[rgb]{0,0,0}\frac{b}{\gamma}}\int_{-\tau}^{0}x_{0}(t-s-\tau)\phi(s)ds$
$\quad+{\color[rgb]{0,0,0}\frac{1}{\gamma}}\int_{0}^{t}x_{0}(t-s)\xi(s)ds,$ (2)
@@
In Appendix B, we show that the fundamental solution can be expanded as
$x_{0}(t)=\sum_{k=-\infty}^{\infty}\dfrac{{\color[rgb]{0,0,0}\gamma}}{{\color[rgb]{0,0,0}\gamma}+b\tau e^{-S_{k}\tau}}e^{S_{k}t}\ \ (t>0),$ (4)
@@
$\langle x(t)\rangle$ $=x_{0}(t)\phi(0)+{\color[rgb]{0,0,0}\frac{b}{\gamma}}\int_{-\tau}^{0}x_{0}(t-s-\tau)\phi(s)ds,$ (6)
$\langle x^{2}(t)\rangle$ $-\langle x(t)\rangle^{2}={\color[rgb]{0,0,0}\frac{2T}{\gamma}}\int_{0}^{t}x_{0}^{2}(s)ds.$ (7)
Note that the history function $\phi(t)$ enters into the expression only for the mean value, but not for the variance. Parameters $a,b,\tau$ in Eq. (5) determine the distribution of characteristic roots $\{S_{k}\}$, which further control the asymptotic behavior of the solution $x(t)$ . If $\max\{\text{Re}[S_{k}]\}<0$, the distribution of $x(t)$ becomes stationary as $t\rightarrow\infty$ and the dynamics is stable; if $\max\{\text{Re}[S_{k}]\}=0$, the variance of $x(t)$ diverges linearly in time and the dynamics is marginally stable; if $\max\{\text{Re}[S_{k}]\}>0$, the dynamics exponentially diverges in time.
@@
Figure 1: (a) A single trajectory simulation of the dynamics with diffusive criticality. Here a=5a=-5, $b=5$, $\tau=1$, $T=1$, $\color[rgb]{0,0,0}\gamma=1$}, and $\phi(t)=0\ \ (t\leq 0)$. (b) 100 trajectories of the dynamics with diffusive criticality for each set of parameters. Blue: $a=0.5$ and b=0.5b=-0.5; red: $a=b=0$ (Brownian motion) ; green: a=1a=-1 and $b=1$. Here $T=1$, $\color[rgb]{0,0,0}\gamma=1$}, $\tau=1$, and $\phi(t)=0\ \ (t\leq 0)$.
@@
$a+b=0,\ a\tau<{\color[rgb]{0,0,0}\gamma},$ (8)
there is exactly one zero root in $\{S_{k}\}$ and all the others have negative real parts (we exclude $a\tau={\color[rgb]{0,0,0}\gamma}$ as the coefficients in Eq. (4) diverge). Such critical dynamics corresponds to scenarios of systems under Pyragas control [51, 41, 42, 43]. Under this condition (Eq. (8)), the fundamental solution at large times $t\gg 1/\beta_{1}$ satisfies
$x_{0}(t)=\dfrac{{\color[rgb]{0,0,0}\gamma}}{{\color[rgb]{0,0,0}\gamma}-a\tau}+O(e^{-\beta_{1}t}).$ (9)
where $\beta_{1}=\underset{\mathrm{Re}[S_{k}]<0}{\mathrm{Inf}}\Big|\mathrm{Re}[S_{k}]\Big|$ for $\{S_{k}\}$ in Eq. (5) (see Appendix C for discussions on asymptotic behavior of the fundamental solution; the same discussion applies to Eq. (14) at below). Combining Eq. (9) with Eqs. (6) and (7), we obtain the behaviors for $t\gg 1/\beta_{1}$:
$\langle x(t)\rangle=\frac{{\color[rgb]{0,0,0}\gamma}}{{\color[rgb]{0,0,0}\gamma}-a\tau}\phi(0)-\frac{a}{{\color[rgb]{0,0,0}\gamma}-a\tau}\int_{-\tau}^{0}\phi(s)ds+{\color[rgb]{0,0,0}O(e^{-\beta_{1}t})},\ \$(10)
$\langle x^{2}(t)\rangle-\langle x(t)\rangle^{2}=\frac{2{\color[rgb]{0,0,0}\gamma}T}{({\color[rgb]{0,0,0}\gamma}-a\tau)^{2}}\cdot t+C_{0}+{\color[rgb]{0,0,0}O(e^{-\beta_{1}t})},\ \$(11)
@@
$C_{0}=\frac{2T}{{\color[rgb]{0,0,0}\gamma}}\int_{0}^{\infty}\left[x_{0}^{2}(s)-\left(\dfrac{{\color[rgb]{0,0,0}\gamma}}{{\color[rgb]{0,0,0}\gamma}-a\tau}\right)^{2}\right]ds,$ (12)
is a bounded constant (see Appendix C for a proof; a similar proof applies to the boundedness of $C_{1}$ in Eq. (18) below). Note that Eq. (11) has been analytically derived in Ref. [52]. The mean displacement $\langle x(t)\rangle$ asymptotically approaches a constant that depends on the historical trajectory $\phi(t)$. The diffusion of the critical dynamics resembles free Brownian motion with a scaling factor $1/({\color[rgb]{0,0,0}\gamma}-a\tau)^{2}$ (Fig. 1(a)). A positive time-delayed force (b=a>0b=-a>0) suppresses the diffusion and a negative one ($b=-a<0,\ a\tau<{\color[rgb]{0,0,0}\gamma}$) enlarges the diffusion (Fig. 1(b)).
@@
there are exactly two conjugate imaginary roots $\pm i\sqrt{b^{2}-a^{2}}/{\color[rgb]{0,0,0}\gamma}$ and all the others have negative real parts. Note that in oscillatory critical dynamics, $\tau$ is uniquely specified by $a$ and $b$. Such oscillatory critical dynamics shows oscillation with a fluctuating and diffusive amplitude (Fig. 2(a)). The fundamental solution in the long-time limit is
$x_{0}(t)$ $=\frac{{\color[rgb]{0,0,0}\gamma}}{{\color[rgb]{0,0,0}\gamma}+b{\color[rgb]{0,0,0}\tau_{c}}e^{-i\omega_{c}{\color[rgb]{0,0,0}\tau_{c}}}}e^{i\omega_{c}t}+\frac{{\color[rgb]{0,0,0}\gamma}}{{\color[rgb]{0,0,0}\gamma}+b{\color[rgb]{0,0,0}\tau_{c}}e^{i\omega_{c}{\color[rgb]{0,0,0}\tau_{c}}}}e^{-i\omega_{c}t}+O(e^{-\beta_{1}t})$
$=\dfrac{2{\color[rgb]{0,0,0}\gamma}\cos(\omega_{c}t+\theta)}{\sqrt{{\color[rgb]{0,0,0}\gamma^{2}}+b^{2}{\color[rgb]{0,0,0}\tau_{c}}^{2}-2{\color[rgb]{0,0,0}\gamma}a{\color[rgb]{0,0,0}\tau_{c}}}}+{\color[rgb]{0,0,0}O(e^{-\beta_{1}t})},$ (14)
@@
$\omega_{c}=\frac{\sqrt{b^{2}-a^{2}}}{{\color[rgb]{0,0,0}\gamma}}\ \ \text{and}\ \ \theta=\arctan\left(\frac{\sqrt{b^{2}-a^{2}}\cdot{\color[rgb]{0,0,0}\tau_{c}}}{a\tau-{\color[rgb]{0,0,0}\gamma}}\right).$ (15)
The mean displacement at large times $t\gg 1/\beta_{1}$ is
$\langle x(t)\rangle=\dfrac{2{\color[rgb]{0,0,0}\gamma}}{\sqrt{{\color[rgb]{0,0,0}\gamma^{2}}+b^{2}{\color[rgb]{0,0,0}\tau_{c}}^{2}-2{\color[rgb]{0,0,0}\gamma}a{\color[rgb]{0,0,0}\tau_{c}}}}$
$\cdot\left(\phi(0)\cos(\omega_{c}t+\theta)+\frac{b}{{\color[rgb]{0,0,0}\gamma}}\int_{-{\color[rgb]{0,0,0}\tau_{c}}}^{0}\cos[\omega_{c}(t-s-{\color[rgb]{0,0,0}\tau_{c}})+\theta]\phi(s)ds\right)$
$\ \ \ \ +{\color[rgb]{0,0,0}O(e^{-\beta_{1}t})},$ (16)
and the variance is
@@
$\dfrac{4{\color[rgb]{0,0,0}\gamma}T\bigg[t+\frac{1}{2\omega_{c}}\bigg(\sin[2(\omega_{c}t+\theta)]-\sin(2\theta)\bigg)\bigg]}{{\color[rgb]{0,0,0}\gamma^{2}}+b^{2}{\color[rgb]{0,0,0}\tau_{c}}^{2}-2{\color[rgb]{0,0,0}\gamma}a{\color[rgb]{0,0,0}\tau_{c}}}+C_{1}+{\color[rgb]{0,0,0}O(e^{-\beta_{1}t})},$ (17)
@@
$C_{1}=\frac{2T}{{\color[rgb]{0,0,0}\gamma}}\int_{0}^{\infty}\bigg(x_{0}^{2}(s)-\dfrac{4{\color[rgb]{0,0,0}\gamma^{2}}\cos^{2}(\omega_{c}s+\theta)}{{\color[rgb]{0,0,0}\gamma^{2}}+b^{2}{\color[rgb]{0,0,0}\tau_{c}}^{2}-2{\color[rgb]{0,0,0}\gamma}a{\color[rgb]{0,0,0}\tau_{c}}}\bigg)ds,$ (18)
@@
Figure 2: (a) A single trajectory simulation of the dynamics with oscillatory criticality. Here a=2a=-2, b=3b=-3, $\tau\approx 1.029$, $T=1$, $\color[rgb]{0,0,0}\gamma=1$}, and $\phi(t)=0\ \ (t\leq 0)$. (b) 100 trajectories of the dynamics with oscillatory criticality for each set of parameters. Blue: a=2a=-2, b=10b=-10, and $\tau\approx 0.181$; red: a=2a=-2, b=5b=-5, and $\tau\approx 0.433$; green: a=2a=-2, b=2.5b=-2.5, and $\tau\approx 1.665$. Here $T=1$, $\color[rgb]{0,0,0}\gamma=1$}, and $\phi(t)=0\ \ (t\leq 0)$.
@@
In the following, we analytically derive the asymptotic expressions for the average heat dissipation rate of two types of marginally stable dynamics, along with numerical verification. This section along with Sec. IV constitute the main results of our work.
@@
where $\circ$ denotes the Stratonovich product [53]. It represents the heat dissipated from the system into the surrounding thermal medium per unit time. A positive sign indicates that the energy is released from the system to the environment. The heat dissipation rate of steady state for stable solutions has been derived in [26, 27, 31]. However, those methods are not applicable when calculating the heat dissipation of marginally stable dynamics. In this section, we explicitly calculate the heat dissipation for the two classes of marginally stable dynamics.
From the definition in Eq. (19) and the dynamics described by Eq. (1), the heat dissipation of the system within an infinitesimal time interval $dt$ for $\tau>0$ is given by ($\gamma\equiv 1$):
$\langle dq(t)\rangle=\left\langle\bigg[{\color[rgb]{0,0,0}\gamma}\frac{dx(t)}{dt}-\xi(t)\bigg]\circ dx(t)\right\rangle$
$=\left\langle\big[ax(t)+bx(t-\tau)\big]\circ{\color[rgb]{0,0,0}\frac{1}{\gamma}}\big[ax(t)+bx(t-\tau)+\xi(t)\big]dt\right\rangle$
$=\frac{a^{2}}{{\color[rgb]{0,0,0}\gamma}}\langle x^{2}(t)\rangle dt+\frac{b^{2}}{{\color[rgb]{0,0,0}\gamma}}\langle x^{2}(t-\tau)\rangle dt$
$\quad\quad+\frac{2ab}{{\color[rgb]{0,0,0}\gamma}}\langle x(t)x(t-\tau)\rangle dt+\frac{aT}{{\color[rgb]{0,0,0}\gamma}}dt,$ (20)
@@
$\langle x(t-\tau)\circ\xi(t)dt\rangle=\left\langle\frac{x(t-\tau+dt)+x(t-\tau-dt)}{2}\cdot\xi(t)dt\right\rangle$
$=\frac{1}{2\gamma}\int_{0}^{t-\tau+dt}x_{0}(t-\tau+dt-s)\langle\xi(s)ds\cdot\xi(t)dt\rangle$
$\quad\quad+\frac{1}{2\gamma}\int_{0}^{t-\tau-dt}x_{0}(t-\tau-dt-s)\langle\xi(s)ds\cdot\xi(t)dt\rangle$
$=\frac{1}{2\gamma}\left(\int_{0}^{t-\tau+dt}x_{0}(t-\tau+dt-s)\cdot 2\gamma T\delta(t-s)ds\right)\cdot dt$
$=\frac{1}{2\gamma}\cdot x_{0}(0)\cdot 2\gamma T\cdot\delta_{\tau}dt=T\delta_{\tau}dt.\quad(\text{since}\ x_{0}(0)=1)$ (21)
Here, $dt$ denotes an infinitesimal time interval and $\delta_{\tau}=1$ when $\tau=0$ and $\delta_{\tau}=0$ for $\tau>0$. For the time-delayed dynamics, the term $\langle bx(t-\tau)\circ\xi(t)\rangle$ equals zero in the calculation of Eq. (III) because the system’s state at $t-\tau$ is uncorrelated with the noise at $t$ [27]. The correlation $\langle x(t)x(t-\tau)\rangle$ can be calculated through
@@
$=\frac{a}{{\color[rgb]{0,0,0}\gamma}}\langle x^{2}(t)\rangle dt+\frac{b}{{\color[rgb]{0,0,0}\gamma}}\langle x(t)x(t-\tau)\rangle dt+{\color[rgb]{0,0,0}\frac{1}{\gamma}}\langle x(t)\circ\xi(t)dt\rangle,$ (22)
@@
$\frac{b}{{\color[rgb]{0,0,0}\gamma}}\langle x(t)x(t-\tau)\rangle=\frac{1}{2}\frac{d\langle x^{2}(t)\rangle}{dt}-\frac{a}{{\color[rgb]{0,0,0}\gamma}}\langle x^{2}(t)\rangle-\frac{T}{{\color[rgb]{0,0,0}\gamma}}.$ (23)
@@
$\left\langle\frac{dq}{dt}\right\rangle=\frac{b^{2}}{{\color[rgb]{0,0,0}\gamma}}\langle x^{2}(t-\tau)\rangle-\frac{a^{2}}{{\color[rgb]{0,0,0}\gamma}}\langle x^{2}(t)\rangle+a\frac{d\langle x^{2}(t)\rangle}{dt}-\frac{aT}{{\color[rgb]{0,0,0}\gamma}}.$ (24)
Note that this expression holds for any $\tau>0$ and $t>0$ in spite of the stability of the dynamics. Below, we discuss the heat dissipation rate of two types of marginally stable dynamics.